# May 28, 2024

### v1

@@ Abstract.
For $n\geq 5$ Fanelli, Floris and Zimmermann gave examples of connected, algebraic subgroups of $\operatorname{Bir}({\mathbb{P}}^{n})$ that are not contained in any maximal one, answering a question of Blanc [BFT23]; see [FFZ23]. The proof relies on the existence of a stably rational but nonrational conic bundle 3-fold $X$ [BCTSSD85], a careful investigation of the birational geometry of unstable ${\mathbb{P}}^{1}$-bundles over $X$, and a study of unipotent group actions. For background and previous results, see [BFT22, BFT23, Fon23] for birational group actions, and [Sar80, Cor95, KM98, KSC04] for minimal models and birational maps.
In this note we follow [Fon23, FFZ23], but, instead of birational geometry, we focus entirely on the action of the unipotent radical of an algebraic subgroup of $\operatorname{Bir}({\mathbb{P}}^{n})$. As a consequence, we show that there are many such non-maximal subgroups, and obtain examples in $\operatorname{Bir}({\mathbb{P}}^{4})$ as well.
Our results apply over algebraically closed fields of arbitrary characteristic.
@@
For Corollaries 24 we use the notation and assumptions of Theorem 1.
@@
The next example shows that the assumption $(X,L)\not\cong\bigl{(}{\mathbb{P}}^{n},{\mathcal{O}}_{{\mathbb{P}}^{n}}(1)\bigr{)}$ is necessary in Theorem 1. (Strictly speaking, ${\mathcal{O}}_{{\mathbb{P}}^{n}}(1)$ does not have a $\mathrm{PGL}_{n+1}$-linearization. But is has an $\mathrm{SL}_{n+1}$-linearization, which is enough for Theorem 1.)
@@ 1. Groups acting on ruled varieties
Let $G\subset\operatorname{Bir}(Y)$ be an algebraic subgroup. If a normal, unipotent subgroup $U\subset G$ acts on $Y$ with small orbits, then every $G\subset\bar{G}\subset\operatorname{Bir}(Y)$ is likely to have a normal, unipotent subgroup $\bar{U}\subset\bar{G}$ that acts on $Y$ with the same (generic) orbits.
@@
Let $\operatorname{Bir}\bigl{(}X\times({\mathbb{P}}^{1},\infty)\to X\bigr{)}\subset\operatorname{Bir}(X\times{\mathbb{P}}^{1})$ denote the subgroup of those birational maps $\phi:X\times{\mathbb{P}}^{1}\dasharrow X\times{\mathbb{P}}^{1}$ that commute with projection to $X$ and map $X\times\{\infty\}$ to $X\times\{\infty\}$. That is, there is a birational map $\phi_{X}:X\dasharrow X$ and a commutative diagram
@@
Let $X$ be a normal, projective variety. Let $G\subset\operatorname{Bir}\bigl{(}X\times({\mathbb{P}}^{1},\infty)\to X\bigr{)}$ be a connected, algebraic subgroup. Assume that
@@
The closures of the $R^{\prime}$-orbits are still the fibers of the projection $\pi:X\times{\mathbb{P}}^{1}\to X$, thus $N_{\bar{G}}(R^{\prime})^{\circ}$ birationally preserves $\pi:X\times{\mathbb{P}}^{1}\to X$. The section $X\times\{\infty\}$ is the fixed point set of the $R^{\prime}$-action, so this is also preserved. Thus
@@ 2. Lemmas on algebraic groups
For an algebraic group $G$, we let $G^{\circ}$ denote the maximal, connected subgroup. $B_{G}\subset G$ denotes a Borel (=maximal, connected, solvable) subgroup, $U_{G}\subset B_{G}$ a maximal, connected, unipotent subgroup, and $R_{G}\subset U_{G}$ the unipotent radical of $G$. $Z(\ )$ denotes the center, and $N_{G}(\ )$ the normalizer in $G$.
@@
Otherwise $U_{G}=U_{G^{\prime}}$. For every simple root of $G^{\prime}$ there is a minimal parabolic subgroup $B_{G^{\prime}}\subsetneq P_{i}\subset G^{\prime}$; let $V_{i}\subset U_{G}$ be its unipotent radical. Then $V_{i}\subset U_{G^{\prime}}=U_{G}$ has codimension 1 and $P_{i}\subset N_{G^{\prime}}(V_{i})$.
@@ 3. Regularizing group actions
Note that the natural way to get a vector bundle out of $P$ is to take $\pi_{*}{\mathcal{O}}_{P}(D)$. This sits in the exact sequence
@@
Let $\bar{X}$ be a normal, projective variety with smooth locus $X\subset\bar{X}$ and $G\subset\operatorname{Bir}\bigl{(}X\times({\mathbb{P}}^{1},\infty)\to X\bigr{)}$ an algebraic subgroup. Assume that $G^{h}\subset\operatorname{Aut}(\bar{X})=\operatorname{Aut}(X)$. Then there is a $G^{h}$-invariant open subset $X^{\circ}\subset X$, and a ${\mathbb{P}}^{1}$-bundle $\pi^{\circ}:P^{\circ}\to X^{\circ}$ with a rational section $D^{\circ}\subset P^{\circ}$, such that
@@
Using the notation and assumptions of (19), assume in addition that there is a ${\mathbb{G}}_{m}\cong T\subset G^{v}/U^{v}_{G}$ and $\dim U^{v}_{G}\geq 1$. Then there is a unique ${\mathbb{P}}^{1}$-bundle $\pi^{m}:P^{m}\to X$ for which
  1. (20.1)
    $D^{m}\subset P^{m}$ is a section,
@@ 4. Proof of Theorem 1
Then $X$ is a cone and the ${\mathbb{G}}_{a}$-action is along the generators.
@@ Acknowledgments.
Partial financial support was provided by the NSF under grant number DMS-1901855.

### v2

# June 4, 2024

@@ Abstract.
Fong and Zikas gave examples of connected, algebraic subgroups of $\operatorname{Bir}(C\times{\mathbb{P}}^{n})$ that are not contained in any maximal one, where $C$ is a non-rational curve [Fon24, FZ23]. Fanelli, Floris and Zimmermann constructed similar subgroups in $\operatorname{Bir}({\mathbb{P}}^{n})$ for $n\geq 5$, answering a question of Blanc [BFT23]; see [FFZ23]. The proof in [FFZ23] relies on the existence of a stably rational but nonrational conic bundle 3-fold $X$ [BCTSSD85], a careful investigation of the birational geometry of unstable ${\mathbb{P}}^{1}$-bundles over $X$, and a study of unipotent group actions. For background and previous results, see [BFT22, BFT23] for birational group actions, and [Sar80, Cor95, KM98, KSC04] for minimal models and birational maps.
In this note we follow [Fon24, FZ23, FFZ23], but, instead of using birational geometry, we focus entirely on the action of the unipotent radical of an algebraic group. This leads to examples in $\operatorname{Bir}({\mathbb{P}}^{4})$ as well.
As another consequence, our results apply over algebraically closed fields of arbitrary characteristic.
@@
Note that [Fon24, FZ23, FFZ23] consider cases for which $\operatorname{Aut}^{\circ}(X)=\{1\}$, and use a Noether-Fano type argument to show that $G$ preserves the negative section of ${\mathbb{P}}_{X}({\mathcal{O}}_{X}{\oplus}L)$. It is quite likely that one can prove Theorem 1 using the methods of [Fon24, FZ23, FFZ23], at least in characteristic 0.
For Corollaries 24 we use the notation and assumptions of Theorem 1, and the description of $\operatorname{Aut}^{\circ}\bigl{(}{\mathbb{P}}_{X}({\mathcal{O}}_{X}{\oplus}L)\bigr{)}$ given in Paragraph 18.
@@
The construction in [FFZ23] uses a non-ample line bundle; these can be treated using Theorem 12.
The next example shows that the assumption $(X,L)\not\cong\bigl{(}{\mathbb{P}}^{n},{\mathcal{O}}_{{\mathbb{P}}^{n}}(1)\bigr{)}$ is necessary in Theorem 1. (Strictly speaking, ${\mathcal{O}}_{{\mathbb{P}}^{n}}(1)$ does not have a $\mathrm{PGL}_{n+1}$-linearization for $n>0$. However, is has an $\mathrm{SL}_{n+1}$-linearization, which is enough for Theorem 1.)
@@ 1. Groups acting on ruled varieties
Let $G\subset\operatorname{Bir}(Y)$ be an algebraic subgroup. If a normal, unipotent subgroup $U\subset G$ acts on $Y$ with small orbits, then every $G\subset\bar{G}\subset\operatorname{Bir}(Y)$ is likely to have a normal, unipotent subgroup $\bar{U}\subset\bar{G}$ that acts on $Y$ with the same orbits.
@@
Let $\operatorname{Bir}\bigl{(}X\times({\mathbb{P}}^{1},\infty)\to X\bigr{)}\subset\operatorname{Bir}(X\times{\mathbb{P}}^{1})$ denote the subgroup of those birational maps $\phi:X\times{\mathbb{P}}^{1}\dasharrow X\times{\mathbb{P}}^{1}$ that commute with projection to $X$ and map $X\times\{\infty\}$ birationally to $X\times\{\infty\}$. Thus there is a birational map $\phi_{X}:X\dasharrow X$ and a commutative diagram
@@
Let $X$ be a normal, projective variety and $G\subset\operatorname{Bir}\bigl{(}X\times({\mathbb{P}}^{1},\infty)\to X\bigr{)}$ a connected, algebraic subgroup. Assume that
@@
A key point is that generically $R^{\prime}$ and $U^{v}_{G}$ have the same orbits. In particular, the closures of the 1-dimensional $R^{\prime}$-orbits are still the fibers of the projection $\pi:X\times{\mathbb{P}}^{1}\to X$. Thus $N_{\bar{G}}(R^{\prime})^{\circ}$ birationally preserves $\pi:X\times{\mathbb{P}}^{1}\to X$. The section $X\times\{\infty\}$ is the fixed point set of the $R^{\prime}$-action, so this is also preserved. Thus
@@ 2. Lemmas on algebraic groups
For an algebraic group $G$, we let $G^{\circ}$ denote the maximal, connected subgroup. $B_{G}\subset G$ denotes a Borel subgroup, $U_{G}\subset B_{G}$ a maximal, connected, unipotent subgroup, and $R_{G}\subset U_{G}$ the unipotent radical of $G$. $Z(\ )$ denotes the center, and $N_{G}(\ )$ the normalizer in $G$.
@@
Otherwise $U_{G}=U_{G^{\prime}}$. For every simple root of $G^{\prime}$ there is a minimal parabolic subgroup $B_{G^{\prime}}\subsetneq P_{i}\subset G^{\prime}$; let $V_{i}\subset U_{G}$ be its unipotent radical. Then $P_{i}/V_{i}$ is either $\mathrm{SL}_{2}$ or $\mathrm{PSL}_{2}$. Therefore $V_{i}\subset U_{G^{\prime}}=U_{G}$ has codimension 1 and $P_{i}\subset N_{G^{\prime}}(V_{i})$.
@@ 3. Regularizing group actions
Note that the natural way to get a vector bundle on $X$ is to take $\pi_{*}{\mathcal{O}}_{P}(D)$. This sits in the exact sequence
@@
Let $\bar{X}$ be a normal, projective variety with smooth locus $X\subset\bar{X}$, and $G\subset\operatorname{Bir}\bigl{(}X\times({\mathbb{P}}^{1},\infty)\to X\bigr{)}$ an algebraic subgroup. Assume that $G^{h}\subset\operatorname{Aut}(\bar{X})=\operatorname{Aut}(X)$. Then there is a $G^{h}$-invariant open subset $X^{\circ}\subset X$, and a ${\mathbb{P}}^{1}$-bundle $\pi^{\circ}:P^{\circ}\to X^{\circ}$ with a rational section $D^{\circ}\subset P^{\circ}$, such that
@@
Using the notation and assumptions of (19), assume in addition that there is a ${\mathbb{G}}_{m}\cong T\subset G^{v}/U^{v}_{G}$ and $\dim U^{v}_{G}\geq 1$. Then there is a unique ${\mathbb{P}}^{1}$-bundle $\pi^{m}:P^{m}\to X$ with a $G$-invariant section $D^{m}\subset P^{m}$, such that
@@ 4. Proof of Theorem 1
Then $X$ is a cone, and the ${\mathbb{G}}_{a}$-action is along its generators.
@@ Acknowledgments.
I thank E. Floris and P. Fong for helpful comments. Partial financial support was provided by the NSF under grant number DMS-1901855.